OpenStudy (magentama202):

WILL FAN AND MEDAL any who help: Choose the correct simplification of (7x^3y^3)^2. Select one: a. 14x^6 y^6 b. 49x^5 y^5 c. 49x^6 y^6 d. 14x^5y^5

OpenStudy (magentama202):

@.Sam. @mjdennis @agent0smith @undeadknight26 @phi @Preetha @mathmate

OpenStudy (magentama202):

I think for this one you would do similar steps according to this formula elsewhere in the lesson (xy)^n = x^ny^n To raise a product to a power, apply that power to each factor inside the parentheses.

OpenStudy (magentama202):

but its also raising a power to a power so it would also be this formula: (x^m)^n = x^(m • n) To raise a power to a power, multiply the exponents.

OpenStudy (agent0smith):

Both formulas tell you the same thing.

OpenStudy (magentama202):

?

OpenStudy (agent0smith):

Essentially they're the same thing. Start attempting to simplify it :)

OpenStudy (magentama202):

ok

OpenStudy (magentama202):

\[(7x^3y^3)^2\] so you would multiply 2 by 3 for 7x^3 and y^3?

OpenStudy (magentama202):

so it would be \[14x^6 y^6\]

OpenStudy (agent0smith):

Remember that the 7 already has a 1 exponent on it.

OpenStudy (magentama202):

so what does that mean?

OpenStudy (agent0smith):

Treat the 7 the same way you did the x and y.

OpenStudy (magentama202):

so .. wait do I add the 1 exponent to the 2?

OpenStudy (magentama202):

*multiply

OpenStudy (agent0smith):

Treat the 1 exponent on the 7, the exact same way you did the exponents on the x and y. Remember, using (x^m)^n = x^(m • n)

OpenStudy (magentama202):

so 1 + 3 * the 2?? 7 (7x^3y^3)^2

OpenStudy (magentama202):

i think im starting to confuse myself on something simple

OpenStudy (agent0smith):

You're still forgetting the 1 exponent on the 7. Literally put one there, \[\large (7^1 x^3y^3)^2\]Now simplify.

OpenStudy (magentama202):

yes thats what i ment by the 1+3 i just forgot to put it in the equation

OpenStudy (magentama202):

oh would you multiply 1*3*2?

OpenStudy (magentama202):

but that would still just be 6 :(

OpenStudy (magentama202):

wait could it be: \[49x^6y^6\]

OpenStudy (magentama202):

7^2 power would be 49

OpenStudy (magentama202):

and the 1*2 would be that 2 i was over thinking it!!

OpenStudy (magentama202):

thanks so much @age

OpenStudy (magentama202):

@agent0smith

OpenStudy (magentama202):

sorry sent the 3rd to last with out your whole name

OpenStudy (agent0smith):

\[\large (7^1 x^3y^3)^2 = 7^2 x^6 y^6\] Good job :) That;s why whenever there is a number in the parentheses, always put a 1 exponent on it.

OpenStudy (magentama202):

Thanks!!

OpenStudy (agent0smith):

You're welcome!

OpenStudy (magentama202):

I'm gonna close this one now